STRUCTURE OF LIQUIDS AND LOW-FREQUENCY RAMAN EFFECT

E. N. SHERMATOV, U.N. RADJABOV, Samarkand State University, Univ. bulv 15, 703004 Samarkand, Uzbekistan.

Intensity distribution in line contour of low-frequency Raman effect up to 250 cm^{−1} from exciting light frequency for stokes and antistokes components in number of organic liqiuds were studied. Contour of line may be present as a central component and two lateral components Obtained results were explained in the context about liquid state structure proposed authors. The maxima lateral components give finds about quantity of spatial by extension. And difference of maxima frequencies for lateral components give finds about value of the change for selforganization extension at oscillations. Line width of central component contains the findings about nonuniform process in media.